1)Find the x and y intercepts of-4x + 4y = -4
x 1 y -3/2
x 9/4 y -3/2
x 9/4 y -1
x1 y -1
2)Find the slope of the line (3,4) (6,5)
1/3
5/6
-1/3
3
Sorry I posted 2 questions but I have a large packet thats due asap and these both have to do with graphs

Found 2 solutions by checkley77, richwmiller:
Answer by checkley77(12569)

-4x + 4y = -4
4y=4x-4
y=4x/4-4/4
y=x-1
Intercepts (0,-1) (1,0)
(3,4) (6,5)
Slope=(y2-y1)/(x2-x1)
Slope=(5-4)/(6-3)
Slope=1/3 ans.

Answer by richwmiller(9135)

Reminder ;
You are limited to 4 questions daily!
one question per submission and
no similar problems
We are not here to do your homework but to help you.
You should be learning from the help you get on the first similar problem so that you can solve the rest of the similar problems on your own.
If you need help with more problems daily, you probably need more help than we can give you. You need a paid tutor to analyze what you are missing!
Sad but realistic!